5/22-23: MZ release from PSL
5/20-24: Exchanging cryo valves.
5/27-6/2: Y arm mass lock. (not done yet)
5/29: RT PC maintenance -> go to safe mode for all VIS connected Dolphin (specially BS)
6/3-6/8: Installation of TMS for ETMY
6/8: Engineering run
6/18 Move PR3
- 6/19 DRMI locked.
- Long term:
4/22-5/31: Yarm commissioning
- 5/27-6/30: DRMI commissioning
- 7/1-8/9: DRFPMI commissioning
- 8/12-10/4: DRFPMI ASC
- 10/7-11/15: Post Commissioning
- 11/18-12/13: Noise hunting
- Please reduce the number or speed of 16kHz channels.
Report of work
- COM: DRMI was locked stably for several minutes.
- Earthquake happened last night. Mirrors were mis-alignment, but they were re-aligned.
- Glitches were much less than day before yesterday since PR2 moved to avoid HOM resonance.
- Type Bp: PR2 was moved ~2mm, instead of PR3.
- IOO, COM: OMC work today afternoon.
- DGS: maintenance for exchanging IMC and one of the SR2,3 or PR2,3.
- DMG, DGS: Another 24TB large volume on E18 was enabled. full data started being written.
- Type Bp: PR3 will be moved today by ~100um, ~2mm
- CRY: ETMY cryo cooler will be restarted because of meta stable.